Awolflover1 (awolflover1):

What is the distance between points (10, 7) and (2, 7) on a coordinate plane?

OpenStudy (anonymous):

\[y = \sqrt{(x_2-x_1)+(y_2-y_1)}\] Distance formula, which is derived from the pythagorean theorem Your points are given as (x1, y1) and (x2,y2) Simply plug them into the formula, or think about it geometrically with the pythagorean theorem.

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 10}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 7}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (ybarrap):

$$ L=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2} $$

OpenStudy (ybarrap):